A student guesses on every question of a​ multiple-choice test that has 9 ​questions, each with 3 possible answers. What is the probability that the student will get at least 7 of the questions​ right?
Answer #1

Solution :

n = 9

p = 0.3333

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X 7) = P(x = 7) + P(x = 8) + P(x = 9)

= (9 C 7) * 0.33337 * (0.6667)2 +  (9 C 8) * 0.33338 * (0.6667)1 +  (9 C 9) * 0.33339 * (0.6667)0

Probability = 0.0083